Home  |  Past Issues  |  Contact Us  | RHIC

 

RHIC Startup

by Fulvia Pilat

Photo of Fulvia PilatThis run is the seventh for RHIC physics operations and over the years we developed a start-up plan that, variations due to different species and energies notwithstanding, has evolved into a reproducible procedure. The blue ring is cooled down to 4K first; while the yellow ring is being cooled, beam activities take place in blue in parallel with PS work following the cold wave in yellow. With both rings cold the set-up phase starts, marked by the development of ramp and collisions. At the end of the set-up phase the beam are turned over to the experiments over night typically with lower luminosity, while beam development activities continue during the day to increase luminosity. This latter phase, or ramp-up phase lasts until luminosity and beam conditions are such to allow the beginning of physics running.

Over the years this general script has been followed despite some detours at times made necessary by unforeseen events. More relevantly, over the years the overall process became faster and more efficient, down from 8 weeks in Run-2 - from having the rings cold to physics - to a record 2 and 1/2 weeks in both Run-5 and Run-6.

The start-up of RHIC operations this year has been challenged on 2 fronts. The first challenge once again has been the uncertainty in program funding, and that as we all know necessitated a considerable delay in the beginning of operations. Implicit in the shifting scheduling are also the conflicting needs to keep the systems constantly ready for beam while utilizing the time for maintenance and reducing power consumption.

The second challenge took the form of an unforeseen problem with the cryogenic system. After a record time cool-down from 80 K (where the rings were held over the summer shut-down) to the operating temperature of 4 K, work with beam in the blue ring started in RHIC on February 19 and during the owl shift we had circulating beam in RHIC. After a few days of beam work at night in blue, unfortunately a warm up of the refrigerator was necessary. The problem in the cryo system was identified as the newly installed heat exchanger HX20, almost certainly caused by turbine oil accumulation at the entrance of the heat exchanger itself. A cryogenic bypass that excludes HX20 from the system was quickly put in place and the cool-down started again. In only 3 days the blue ring was at operating temperature and ready for beam again. The cool-down for yellow is also proceeding very well, with the expectation of having yellow at 4 K by Sunday March 11. Beam work restarted in the RHIC blue ring last night with a good start: the first bunch injected into the machine after a couple of weeks just went straight to circulating, hopefully a good omen for the upcoming set-up and ramp-up phase!

The next priority for RHIC is ramp development and operational use of tune and coupling feedback on the ramp, which can potentially make ramp tuning much faster than the traditional feed forwarding of corrections from one ramp to the next. The priority for the injectors, after the successful commissioning of longitudinal bunch merging in the AGS and of more efficient stripping foils in BtA, is development of bunch intensity for RHIC, with a very promising bunch intensity of ~1.2e9 already recorded at AGS extraction. We should also remember that the injectors will run concurrently with the RHIC run the NSRL program from Booster, BLIP from Linac, and polarized proton development in the AGS. If all goes well we expect to be ready for RHIC physics running on or about April 1st.